A Toyota manufacturing plant in South Carolina (U.S.) imports $15,000 of inputs from Mexico and $10,000 of inputs from Canada and produces $80,000 in vehicles in 2014. From these, Toyota sells $25,000 to consumers in 2014 and $25,000 to firms in 2014. They export $20,000 of these vehicles to Canada and the rest remain unsold at the end of 2014.

What is the contribution/change to GDP in the U.S. from the above information? (enter a number only, no symbols such as $, +, -, etc.)
Answer #1

As per the given data,

Increase in imports = 15000 + 10000 = $25000

Increase in consumption = $25000

Increase in investment = $25000

Increase in exports = $20000

Increase in Inventory of finished goods as investment = $10000

Hence,

Contribution to GDP = 25000 + (25000 + 10000) + (20000-25000)

Contribution to GDP = 55000

So, GDP will increase by 55000.